# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

### v1

@@ II-C2 Quantum Differential Privacy
QDP lifts the notion of DP to the space of quantum states, with the neighboring relation typically defined either in terms of closeness in trace distance [8] or reachability by a single local operation [9]. [^1] Given two quantum states $\rho$ and $\sigma$ of $n$ registers each, call them neighbors if it is possible to reach either $\sigma$ from $\rho$ or $\rho$ from $\sigma$ by performing a general quantum channel on a single register only. We denote two states being neighbors by $\rho\sim\sigma$.
@@ III-A Framework
Set $\mathcal{Q}$ of discriminative pairs: This is a subset of $\mathcal{S}\times\mathcal{S}$ that specifies which pairs of elements from $\mathcal{S}$ should be indistinguishable. Namely, if $(\mathcal{T}_{1},\mathcal{T}_{2})\in\mathcal{Q}$, then the goal of the privacy mechanism is to conceal whether the input belongs to $\mathcal{T}_{1}$ or $\mathcal{T}_{2}$. Here, note that $\rho\in\mathcal{T}_{1}\Rightarrow\rho\notin\mathcal{T}_{2}$. Note that $(\mathcal{T}_{i},\mathcal{T}_{j})\in\mathcal{Q}$ if and only if $(\mathcal{T}_{j},\mathcal{T}_{i})\in\mathcal{Q}$. Proceeding with the same example, we can set
@@
Set $\Theta$ of data distributions: A collection of probability distributions $P_{X}\in\mathcal{P}(\mathcal{X})$ over a finite space $\mathcal{X}$ that indexes an ensemble of density operators $\{\rho^{x}\}_{x\in\mathcal{X}}$. Taking $X\sim P_{X}\in\Theta$, the matrix-valued random variable $\rho^{X}$ models a density operator that is randomly chosen according to $P_{X}$. The set $\Theta$ can be understood as capturing beliefs that the adversary has regarding the state of the system.
@@
QPP allows system designers to explicitly encode their assumptions into the privacy framework. Setting the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ to accurately reflect the application of interest is crucial for obtaining meaningful privacy guarantees and to optimize utility. Explicit assumptions can also help account for ethical or fairness concerns associated with quantum systems; cf. Remark 17 for a concrete example concerning quantum fairness and how it is incorporated within QPP.
@@
is a set of pairs of bipartite states. We then say that $\mathcal{A}$ is $(\varepsilon,\delta)$-QPP with respect to that framework if for all $P_{X}\in\Theta$, Msuperscriptnormal-$\mathrm{M}^{\prime}\in\mathcal{M}^{\prime}$, and $(\omega_{RA}^{\mathcal{R}},\omega_{RA}^{\mathcal{T}})\in\mathcal{G}$, we have
$\mathrm{Tr}\!\left[\mathrm{M}^{\prime}(\mathcal{I}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{R}})\right]\leq e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}^{\prime}(\mathcal{I}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{T}}))\right]+\delta,$ (28)
with $\rho^{\mathcal{R}}$ and $\rho^{\mathcal{T}}$ defined similar to Definition 4.
However, it is unclear whether such a stronger privacy notion would be useful in practical applications. For example, consider $\sigma_{1}\coloneqq|0\rangle\!\langle 0|\otimes\rho^{\mathcal{R}}$ and $\sigma_{2}\coloneqq|1\rangle\!\langle 1|\otimes\rho^{\mathcal{T}}$ with $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$. If a measurement on the reference system can be applied, then a computational-basis measurement distinguishes $\sigma_{1}$ and $\sigma_{2}$ perfectly. Thus, it is important to choose $\mathcal{A}$ appropriately with a practically applicable $\mathcal{M}^{\prime}$, such that the required indistinguishability is achieved. We shall revisit a variant of this framework with quantum divergences in Section IX-B.
@@ III-B Equivalent Formulation of QPP with DL Divergence
the inequality in (30) is equivalent to
@@
which implies that $\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\nu}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]\leq\delta$. Recalling that $\nu\leq\varepsilon$ and noting that $\lambda\mapsto\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\lambda}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]$ is a monotonically decreasing function (cf. [14, Lemma 4.2]), we have
@@ III-C1 Quantum DP
In QDP (Definition 3), secrets are singletons, discriminative pairs comprise states satisfying a neighboring relation, while the measurement class $\mathcal{M}$ includes all possible measurements. QPP recovers the QDP setting by making the following choices:
@@
$\Theta$ $=\mathcal{P}(\mathcal{X}),$
@@
More generally, one may add flexibility to the QDP formulation by considering other subsets $\Theta$. This can be used, for instance, to treat situations in which only certain neighboring pairs are of interest, namely, by choosing the distributions that assign positive probabilities only to those selected density operators. This can be interpreted as adding domain knowledge to the original QDP framework.
@@ III-C2 Quantum Local DP
In quantum local DP (QLDP) [10], we choose secret pairs to be pairs of arbitrary distinct states, while the measurement class includes all possible measurements. Thus, QLDP realizes the same $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ framework as QDP, except that $\mathcal{Q}=\{(\rho,\sigma):\rho,\sigma\in\mathcal{D}\}$ for QLDP.
@@ III-C3 Classical PP
Consider a classical PP framework $(\mathcal{S}_{c},\mathcal{Q}_{c},\Theta_{c})$, as specified in Definition 2. Let the encoding of the database $x\in\mathcal{X}^{n\times k}$ be $\rho^{x}\coloneqq|x\rangle\!\langle x|$, and denote a projective measurement operator corresponding to outcome $y$ as $|y\rangle\!\langle y|$. Here note that $\{|x\rangle\}_{x\in\mathcal{X}}$ and $\{|y\rangle\}_{y\in\mathcal{Y}}$ are respective orthonormal bases formed related to the input and output alphabets of the classical PP mechanism $\mathcal{A}_{c}$. Then classical PP is obtained from QPP by setting
@@
$\mathcal{A}(\rho^{x})=\sum_{y,x^{\prime}}p(y|x)|y\rangle\!\langle x^{\prime}|\rho^{x}|x^{\prime}\rangle\!\langle y|$ (41)
@@ IV Datta–Leditzky Information Spectrum Divergence
We now focus on the DL divergence [14], whose operational interpretation in terms of QPP was provided in the previous section, and we study structural properties thereof, which will be useful when analyzing the QPP framework. We first formulate a primal and dual SDP to compute the DL divergence and then use that to prove joint-quasi convexity, the data-processing inequality under positive, trace-preserving maps, and connections to the smooth max-relative entropy.
@@ IV-A SDP Formulations
The dual forms of these optimization problems are derived from the canonical primal and dual formulations of SDPs, which are respectively given by (cf. [27, Definition 2.20])
@@
$\operatorname{Tr}[X\Phi^{{\dagger}}(Y)]$ $=\operatorname{Tr}\!\left[\begin{bmatrix}\mu&0\\ 0&W\end{bmatrix}\begin{bmatrix}-\operatorname{Tr}[Z]&0\\ 0&Z+\lambda\sigma\end{bmatrix}\right]$ (64)
@@ IV-B Properties
We derive several properties of the DL divergence from (13b), which are subsequently used in the analysis of the QPP framework. Basic properties of the DL divergence, including the data-processing inequality, have been proven in [14, Proposition 4.3]. Here, we generalize the data-processing inequality to hold for arbitrary positive, trace non-increasing maps (beyond the set of quantum channels) and also establish joint-quasi convexity of the DL divergence, along with its connection to the smooth max-relative entropy. The proofs of these properties rely on the SDP formulation from Lemma 1.
@@
Property 1: The statement was proven in [14, Proposition 4.3] for a quantum channel $\mathcal{N}$ by using the inequality
$\mathrm{Tr}\!\left[(\rho-e^{\lambda}\sigma)_{+}\right]\geq\mathrm{Tr}\!\left[\!\left(\mathcal{N}(\rho)-e^{\lambda}\mathcal{N}(\sigma)\right)_{+}\right],$ (84)
which holds for all $\lambda\in\mathbb{R}$. Here, we prove the data-processing inequality, but we generalize it to hold for a positive, trace non-increasing map $\mathcal{N}$. Our derivation relies on the SDP formulation of the DL divergence from (42a).
@@
Property 3: From [38, Appendix B], we have that
@@
The proof is concluded by invoking the following lemma (proven in Appendix A).
@@
Property 4: This follows by invoking Property 3 and using the fact that the smooth max-relative entropy satisfies subadditivity (Appendix B) with

### v2

@@ II-C2 Quantum Differential Privacy
QDP lifts the notion of DP to the space of quantum states, with the neighboring relation typically defined either in terms of closeness in trace distance [13], reachability by a single local operation [14], [^1] Given two quantum states $\rho$ and $\sigma$ of $n$ registers each, call them neighbors if it is possible to reach either $\sigma$ from $\rho$ or $\rho$ from $\sigma$ by performing a general quantum channel on a single register only. or by quantum Wasserstein distance of order 1 [53]. We denote two states being neighbors by $\rho\sim\sigma$.
@@ III-A Framework
Set $\mathcal{Q}$ of discriminative pairs: This is a subset of $\mathcal{S}\times\mathcal{S}$ that specifies which pairs of elements from $\mathcal{S}$ should be indistinguishable. Namely, if $(\mathcal{T}_{1},\mathcal{T}_{2})\in\mathcal{Q}$, then the goal of the privacy mechanism is to conceal whether the input belongs to $\mathcal{T}_{1}$ or $\mathcal{T}_{2}$. Note that $\rho\in\mathcal{T}_{1}\Rightarrow\rho\notin\mathcal{T}_{2}$. We require that $\mathcal{Q}$ is symmetric, i.e., that $(\mathcal{T}_{i},\mathcal{T}_{j})\in\mathcal{Q}$ if and only if $(\mathcal{T}_{j},\mathcal{T}_{i})\in\mathcal{Q}$. Proceeding with the same example, we can set
@@
Set $\Theta$ of data distributions: A collection of probability distributions $P_{X}\in\mathcal{P}(\mathcal{X})$ over a finite space $\mathcal{X}$ that indexes an ensemble of density operators $\{\rho^{x}\}_{x\in\mathcal{X}}$. Taking $X\sim P_{X}\in\Theta$, the matrix-valued random variable $\rho^{X}$ models a density operator that is randomly chosen according to $P_{X}$. Proceeding with the same example, $\{\rho^{x}\}_{x\in\mathcal{X}}=\{\sigma:\sigma\in\mathcal{T}_{i},\ \mathcal{T}_{i}\in\mathcal{S}\}\subset\mathcal{D}(\mathcal{H})$. The set $\Theta$ can be understood as capturing beliefs that the adversary has regarding the state of the system.
In the above example, we have considered a subset of density operators (i.e., $\{\rho^{x}\}_{x\in\mathcal{X}}\subset\mathcal{D}(\mathcal{H})$). There could be applications where we have to consider all density operators. To incorporate this, we choose the following: Fix $k\in\mathbb{N}$ and let $\mathfrak{F}_{k}\subset 2^{\mathcal{D}(\mathcal{H})}$ be the collection of all finite subsets of $\mathcal{D}(\mathcal{H})$ with $k$ elements. For each $\mathcal{F}\in\mathfrak{F}_{k}$, we write $\mathcal{P}(\mathcal{F})$ for the class of all distributions supported on $\mathcal{F}$, and define
$\mathcal{P}_{k}\big{(}\mathcal{D}(\mathcal{H})\big{)}\coloneqq\bigcup_{\mathcal{F}\in\mathfrak{F}_{k}}\mathcal{P}(\mathcal{F}).$ (22)
Every distribution $P\in\mathcal{P}_{k}\big{(}\mathcal{D}(\mathcal{H})\big{)}$ is supported on exactly $k$ density operators. Note that all density operators outside of the underlying finite set comprise of the null set. We associate a random variable $X\sim P=P_{X}$ with each such distribution and write $\mathcal{X}=\mathrm{supp}(P_{X})$ for its support. Note the slight abuse in notation, as the support of $P_{X}$ changes with the distribution, which is not reflected in the generic indexing set $\mathcal{X}$. The set of data distributions in the QPP framework is now taken as $\Theta\subseteq\mathcal{P}_{k}\big{(}\mathcal{D}(\mathcal{H})\big{)}$ for some $k\in\mathbb{N}$.
@@
QPP allows system designers to explicitly encode their assumptions into the privacy framework. Setting the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ to accurately reflect the application of interest is crucial for obtaining meaningful privacy guarantees and to optimize utility. Explicit assumptions can also help account for ethical or fairness concerns associated with quantum systems; cf. Proposition 11 for a concrete example concerning quantum fairness and how it is incorporated within QPP.
@@
is a set of pairs of bipartite states with $\rho^{\mathcal{R}}$ and $\rho^{\mathcal{T}}$ defined similar to Definition 4. We then say that $\mathcal{A}$ is $(\varepsilon,\delta)$-QPP in that framework if for all $P_{X}\in\Theta$, $\mathrm{M}^{\prime}\in\mathcal{M}^{\prime}$, and $(\omega_{RA}^{\mathcal{R}},\omega_{RA}^{\mathcal{T}})\in\mathcal{G}$, we have
$\mathrm{Tr}\!\left[\mathrm{M}^{\prime}(\mathcal{I}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{R}})\right]\leq e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}^{\prime}(\mathcal{I}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{T}}))\right]+\delta.$ (28)
However, it is unclear whether such a stronger privacy notion would be useful in practical applications. For example, consider $\sigma_{1}\coloneqq|0\rangle\!\langle 0|\otimes\rho^{\mathcal{R}}$ and $\sigma_{2}\coloneqq|1\rangle\!\langle 1|\otimes\rho^{\mathcal{T}}$ with $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$. If a measurement on the reference system can be applied, then a computational-basis measurement distinguishes $\sigma_{1}$ and $\sigma_{2}$ perfectly. Thus, it is important to choose $\mathcal{A}$ appropriately with a practically applicable $\mathcal{M}^{\prime}$, such that the required indistinguishability is achieved.
We shall revisit a variant of this framework with quantum divergences in Section IX-B. The strength of the privacy framework is determined by the underlying quantum divergence. However, note that the problems discussed previously are not completely solved by the variant proposed therein.
@@ III-B Equivalent Formulation of QPP with DL Divergence
as a consequence of, e.g., [15, Lemma II.1], the inequality in (30) is equivalent to
@@
which implies that $\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\nu}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]\leq\delta$. Recalling that $\nu\leq\varepsilon$ and noting that $\lambda\mapsto\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\lambda}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]$ is a monotonically decreasing function (cf. [24, Lemma 4.2]), we have
@@
In the following remark, we highlight how the DL divergence also provides a novel characterization for classical PP.
Remark 4 (Classical PP through DL divergence).
For discrete probability distributions $p,q\in\mathcal{P}(\mathcal{Y})$, the DL divergence in Eq. 12b reduces to
$\overline{\mathsf{D}}^{\delta}_{c}(p\|q)\coloneqq\\ \ln\inf\left\{\lambda\geq 0:\sum_{y\in\mathcal{Y}}\max\!\left\{p(y)-\lambda\,q(y),0\right\}\leq\delta\right\}.$ (38)
A randomized mechanism ${A}:\mathcal{X}^{n\times k}\to\mathcal{Y}$ is $(\varepsilon,\delta)$-classical PP in the framework $(\mathcal{S}_{c},\mathcal{Q}_{c},\Theta_{c})$ if for all $P_{X}\in\Theta_{c}$, $(\mathcal{R},\mathcal{T})\in\mathcal{Q}_{c}$ with $P_{X}(\mathcal{R}),P_{X}(\mathcal{T})>0$,
$\overline{\mathsf{D}}^{\delta}_{c}\!\left(P_{A(X)|\mathcal{R}}\|P_{A(X)|\mathcal{T}}\right)\leq\varepsilon,$ (39)
where $P_{A(X)|\mathcal{R}},P_{A(X)|\mathcal{T}}$ are the output distributions conditioned on the secret events $\mathcal{R}$ and $\mathcal{T}$, respectively. See also Remark 8 for further connections to information-theoretic quantities characterizing classical privacy frameworks.
We further note that Lemma 1 below provides a semi-definite programming characterization of the DL divergence, which reduces to a linear program in the classical case.
@@ III-C1 Quantum DP
In QDP (Definition 3), secrets are singletons, discriminative pairs comprise states satisfying a neighboring relation, while the measurement class $\mathcal{M}$ includes all possible measurements. QPP recovers the QDP setting by making the following choices while recalling (22)  [^2] For each pair of states $(\rho,\sigma)$, there exists at least one probability distribution that assigns positive probability for these two states, which recovers the definitions of QDP. :
@@
$\Theta$ $={\mathcal{P}_{2}\big{(}\mathcal{D}(\mathcal{H})\big{)}},$
@@
More generally, one may add flexibility to the QDP formulation by considering other subsets $\Theta$ (i.e., $\Theta\subset\mathcal{P}_{2}\big{(}\mathcal{D}(\mathcal{H})\big{)}$). This can be used, for instance, to treat situations in which only certain neighboring pairs are of interest, namely, by choosing the distributions that assign positive probabilities only to those selected density operators. This can be interpreted as adding domain knowledge to the original QDP framework.
@@ III-C2 Quantum Local DP
In quantum local DP (QLDP) [15]  [^3] QLDP is also known as Local differential privacy (under the ‘extreme setting’, as compared to standard QDP) in Section V.A of [15]. , we choose secret pairs to be pairs of arbitrary distinct states, while the measurement class includes all possible measurements. Thus, QLDP realizes the same $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ framework as QDP, except that $\mathcal{Q}=\{(\rho,\sigma):\rho,\sigma\in\mathcal{D}\}$ for QLDP.
@@ III-C3 Classical PP
Consider a classical PP framework $(\mathcal{S}_{c},\mathcal{Q}_{c},\Theta_{c})$, as specified in Definition 2. Assume that $p_{X}\in\Theta_{c}$ are discrete probability distributions over the probability space $\mathcal{P}(\mathcal{X}^{n\times k})$. Let the encoding of the database $x\in\mathcal{X}^{n\times k}$ be $\rho^{x}\coloneqq|x\rangle\!\langle x|$, and denote a projective measurement operator corresponding to outcome $y$ as $|y\rangle\!\langle y|$. Here note that $\{|x\rangle\}_{x\in\mathcal{X}}$ and $\{|y\rangle\}_{y\in\mathcal{Y}}$ are respective orthonormal bases formed related to the input and output alphabets of the classical PP mechanism $\mathcal{A}_{c}$. Then classical PP is obtained from QPP by setting
@@
$\mathcal{A}(\rho^{x})=\sum_{y\in\mathcal{Y},x^{\prime}\in\mathcal{X}}p(y|x)|y\rangle\!\langle x^{\prime}|\rho^{x}|x^{\prime}\rangle\!\langle y|$ (43)
@@ IV Datta–Leditzky Information Spectrum Divergence
We now focus on the DL divergence [24], whose operational interpretation in terms of QPP was provided in the previous section (see Remark 5), and we study structural properties thereof, which will be useful when analyzing the QPP framework. We first formulate a primal and dual SDP to compute the DL divergence and then use that to prove joint-quasi convexity, the data-processing inequality under positive, trace-preserving maps, and connections to the smooth max-relative entropy.
@@ IV-A SDP Formulations
The dual forms of these optimization problems are derived from the canonical primal and dual formulations of SDPs, which are respectively given by (cf. [37, Definition 2.20])
@@
$\operatorname{Tr}[X\Phi^{{\dagger}}(Y)]$
$=\operatorname{Tr}\!\left[\begin{bmatrix}\mu&0\\ 0&W\end{bmatrix}\begin{bmatrix}-\operatorname{Tr}[Z]&0\\ 0&Z+\lambda\sigma\end{bmatrix}\right]$ (54)
@@
Corollary 1 (Another formulation of the DL divergence).
DL divergence has the following equivalent formulation:
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)=\ln\sup_{0\leq W\leq I,\mathrm{Tr}[W\rho]\geq\delta}\frac{\mathrm{Tr}[W\rho]-\delta}{\mathrm{Tr}[W\sigma]}.$ (61)
Proof:
Consider the SDP formulation in (44d) and set $W^{\prime}=\frac{W}{\mu}$ therein to arrive at
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $=\ln\sup_{\mu,W^{\prime}\geq 0}\left\{\begin{array}[c]{c}\mu\operatorname{Tr}[W^{\prime}\rho]-\mu\delta:\\ \mu\operatorname{Tr}[W^{\prime}\sigma]\leq 1,\ W^{\prime}\leq I\end{array}\right\}$ (64)
$=\ln\sup_{0\leq W^{\prime}\leq I,\mathrm{Tr}[W^{\prime}\rho]\geq\delta}\frac{\mathrm{Tr}[W^{\prime}\rho]-\delta}{\mathrm{Tr}[W^{\prime}\sigma]},$ (65)
where the last equality follows from identifying that $\mu=1/\operatorname{Tr}[W^{\prime}\sigma]$ is the $\mu$ that maximizes the former, given that $\mathrm{Tr}[W^{\prime}\rho]\geq\delta$. When $\mathrm{Tr}[W^{\prime}\rho]<\delta$, the optimum $\mu=0$ and the objective within the supremum becomes zero. Replacing $W^{\prime}$ by $W$, concludes the proof. ∎
Remark 8 (Approximate-max divergence).
In [2], the $\delta$-approximate-max divergence is defined as
$\mathsf{D}_{\infty}^{\delta}(p_{Y}\|p_{Z})\coloneqq\ln\max_{S\in\mathrm{Supp}(Y),\Pr[Y\in S]\geq\delta}\frac{\Pr[Y\in S]-\delta}{\Pr[Z\in S]},$ (66)
where $Y$ and $Z$ are random variables distributed according to $Y\sim p_{Y}$ and $Z\sim p_{Z}$. By substituting classical distributions into Corollary 1, we observe that the DL divergence reduces to the approximate-max divergence. Note that the approximate-max divergence has been used to characterize $(\varepsilon,\delta)$-(classical) DP in [2, Remark 3.1]. Thus, this showcases that the equivalence we established for QPP with the DL divergence herein reduces to the existing equivalence for (classical) DP.
@@ IV-B Properties
We derive several properties of the DL divergence from (12b), which are subsequently used in the analysis of the QPP framework. Basic properties of the DL divergence, including the data-processing inequality, have been proven in [24, Proposition 4.3]. Here, we generalize the data-processing inequality to hold for arbitrary positive, trace non-increasing maps (beyond the set of quantum channels) and also establish joint-quasi convexity of the DL divergence, along with its connection to the smooth max-relative entropy (recall the definition in (15)). The proofs of these properties rely on the SDP formulation from Lemma 1.
@@
Property 1: The statement was proven in [24, Proposition 4.3] for a quantum channel $\mathcal{N}$ by using the inequality
$\mathrm{Tr}\!\left[(\rho-e^{\gamma}\sigma)_{+}\right]\geq\mathrm{Tr}\!\left[\left(\mathcal{N}(\rho)-e^{\gamma}\mathcal{N}(\sigma)\right)_{+}\right],$ (76)
which holds for all $\gamma\in\mathbb{R}$. Here, we prove the data-processing inequality, but we generalize it to hold for a positive, trace non-increasing map $\mathcal{N}$. Our derivation relies on the SDP formulation of the DL divergence from (44a).
@@
Note that this property can also be derived using the proof of [38, Lemma 4].
@@
We also present an alternative proof for this using the joint-convexity of hockey-stick divergence in Appendix A.
Property 3: From [48, Appendix B], we have that
@@
The proof is concluded by invoking the following lemma (proven in Appendix B).
@@
Property 4: This follows by invoking Property 3 and using the fact that the smooth max-relative entropy satisfies subadditivity (Appendix C) with